OpenStudy (anonymous):

Which of the following best defines 3 2/3= ? Square root of 9 Cube root of 9 Cube root of 3 Square root of 3

OpenStudy (mathstudent55):

Is there a symbol missing somewhere?

OpenStudy (mathstudent55):

Can you draw it?

OpenStudy (♪chibiterasu):

I think she means: \[3 \frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

Yes that's what I mean

OpenStudy (mathmale):

\[3 \frac{ 2 }{ 3 }\] simplifies to what? 2 is the cube root of what number?

OpenStudy (mathstudent55):

I don't understand the question, then. Square root of 9 is 3, a rational number. All other choices are irrational numbers. 3 2/3 is a rational number.

OpenStudy (anonymous):

I don't understand the question either, but that's how it's worded.

OpenStudy (mathstudent55):

Given: \(3\dfrac{2}{3} \approx 3.67 \) Choices: A. \(\sqrt{9} = 3\) B. \(\sqrt[3]{9} \approx 2.08\) C. \(\sqrt[3]{3} \approx 1.44\) D. \(\sqrt{3} \approx 1.732\) As you can see, 3 2/3 is closest to A, but I still think this is a strange problem.

OpenStudy (anonymous):

Thank you haha

OpenStudy (mathstudent55):

You're welcome.